Write an explicit formula for the nth term of each arithmetic sequence
6,8,10,12,14.....

Answer by stanbon(75887)

Write an explicit formula for the nth term of each arithmetic sequence
6,8,10,12,14.....
a(n)=a(1)+(n-1)d
In your case: a(1)=6 and d=10-8=2
So, a(n)=6+(n-1)2
Cheers,
Stan H.